# Poisson approximation of prime divisors of shifted primes

### v3

@@ 1. Introduction
The prime factorization of a random integer $n\leqslant x$ is well-approximated, in a probabilistic sense, by a sequence of independent random variables. This approximation is known as the Kubilius model after the pioneering work of Kubilius [30]; see also [7] for a development of the model. In this note we develop an analogous model for the prime factorization of a shifted prime $p+a$ with $a$ fixed and $p$ a prime drawn at random from $(|a|,x]$. We prove a total variation distance estimate for the difference between the model and actual prime factors of shifted primes, and apply it to show that the prime factors of shifted primes in disjoint sets behave like independent Poisson variables. This in turn is used to establish a general transference principle, which states roughly that any property of the prime factors of a random integer $n\leqslant x$, which is weakly dependent on the smallest and largest prime factors of $n$, also holds for the prime factors of a random shifted prime $p+a$ with $p\leqslant x$. This excludes properties such as bounding the size of the largest prime factor of $n$ (a notoriously difficult problem) and the property “$2|n$”, which must be excluded since it occurs for half of integers $n\leqslant x$, but for $\pi(x)-1$ primes $p\leqslant x$ with $n=p+a$ when $a$ is odd, and for only one prime $p$ when $a$ is even.
@@ 1.2. A Kubilius type model for shifted primes
the implied constant in the $\ll$-≪ -symbol depending only on $a,A$ and $\alpha$.
@@
The distribution of the large prime factors of shifted primes is not well understood at present. It is known that for infinitely many primes $p$, $P^{+}(p+a)\leqslant p^{0.2844}$ [31] and for infinitely many primes $p$, $P^{+}(p+a)\geqslant p^{0.677}$ [3]. It is conjectured that analogous statements hold with $0.2844$ replaced by any positive number, and with $0.677$ replaced by any number $<1$. We do not address these problems in this paper.
@@ 1.4. An application
for all but $o(x)$ integers up to $x$, as $x\to\infty$. Denote by $q_{1},\ldots,q_{k}$ the primes in $(\xi(x),y]$, let $T_{1}$ be the set of primes in $[2,\xi(x)]$, let $T_{j}=\{q_{j-1}\}$ for $12\leqslant j\leqslant k+$ 1 and set $f_{i}=\omega(n,T_{i})$ for each $i$. The condition (1.8) is then equivalent to $(\omega(n,T_{1}),\ldots,\omega(n,T_{k+1}))\in\mathscr{R}$ for some set $\mathscr{R}\subseteq{\mathbb{N}}_{0}^{k+1}$. Apply Theorem 4 with $\alpha=\frac{1}{2}$ (allowed since Hypothesis $Z(\gamma)$ holds for some $\gamma>\frac{1}{2}$), $A=1$ and $u=\log_{3}x$, and we see that
@@ 3. Sieving by small primes
[…]
@@
Since #$\#\{b\in\mathcal{A}:q|b\}=\pi(x;qm,-a)+O_{a}(1)$ for primes $q$, heuristically we expect that $\Phi_{m}(x,y)$ is approximately $\pi(x)g_{y}(m)$.
@@
[…]
When $g_{y}(m)\neq 0$, the lemma follows quickly from the “Fundamental Lemma” of the sieve, e.g., Theorem 2.5’ in [27]. For any $D\geqslant y$ we have
(3.5) $\Phi_{m}(x,y)=\pi(x)g_{y}(m)\Big{(}1+O_{\varepsilon}(e^{-(1-\varepsilon)v\log v})\Big{)}+O_{a}\bigg{(}D+\sum_{\begin{subarray}{c}d\leqslant D\\ P^{+}(d)\leqslant y\\ (dm,a)=1\end{subarray}}\Big{|}\pi(x;dm,-a)-\frac{\pi(x)}{\phi(dm)}\Big{|}\bigg{)}$
where
$v=\frac{\log D}{\log y}.$
We take $D=\max(y,x^{\theta}/m)$, so that $v=\max(1,{u}_{m}^{\phantom{2}})$. Thus, $v\log v={u}_{m}^{\phantom{2}}\log({u}_{m}^{\phantom{2}}+1)+O(1)$. Since $m\not\in\mathcal{E}(x,y;\theta,A+1)$, (3.4) implies that
@@
[…]
@@ 4. The proof of Theorem 1
$\mathcal{M}_{1}$ =$=\{m<x^{\theta},m\not\in\mathcal{E}(x,y;\theta,A+1)\},$
$\mathcal{M}_{2}$ =$=\{m<x^{\theta},m\in\mathcal{E}(x,y;\theta,A+1)\},$
$\mathcal{M}_{3}$ =$=\{x^{\theta}\leqslant m\leqslant(x+a)/y\},$
$\mathcal{M}_{4}$ =$=\{(x+a)/y<m\leqslant x+a\},$
$\mathcal{M}_{5}$ =$=\{m>x+a\}.$
Throughout our proof, implied constants may depend on $a$, $A$, $\varepsilon$ and $\theta$ only. Now
@@
where $\qquad\text{where }B_{\ell}=(\theta u-\ell)\log(\theta u-\ell+1)+\ell\log(\ell+1).$
@@
When $m>(x+a)/y$, $\Phi_{m}(x,y)$ is 1 if $m-a$ is prime larger than $|a|+1$ and $m-a\leqslant x$, and zero otherwise. Start with the triangle inequality and (3.4), and use the same analysis as in the case $m\in\mathcal{M}_{3}$:
@@
where we applied (3.2) in the last step, assuming that $\delta$ is sufficiently small. By Theorem 1 of [37] [^1] This is stated only for $a=-1$ but the proof works for any $a$ with no essential changes. and standard bounds for the Dickman function $\rho(u)$ (see, e.g., [40, Ch. III.5]),
#$\#\{|a|+1<p\leqslant x:P^{+}(p+a)\leqslant y\}\ll\pi(x)u\rho(u)\ll\pi(x)e^{-(1-\varepsilon)u\log u}.$

### v4

@@ 1. Introduction
The prime factorization of a random integer $n\leqslant x$ is well-approximated, in a probabilistic sense, by a sequence of independent random variables. This approximation is known as the Kubilius model after the pioneering work of Kubilius [30]; see also [7] for a development of the model. In this note we develop an analogous model for the prime factorization of a shifted prime $p+a$ with $a$ fixed and $p$ a prime drawn at random from $(|a|,x]$. We prove a total variation distance estimate for the difference between the model and actual prime factors of shifted primes, and apply it to show that the prime factors of shifted primes in disjoint sets behave like independent Poisson variables. This in turn is used to establish a general transference principle, which states roughly that any property of the prime factors of a random integer $n\leqslant x$, which is weakly dependent on the smallest and largest prime factors of $n$, also holds for the prime factors of a random shifted prime $p+a$ with $p\leqslant x$. This excludes properties such as bounding the size of the largest prime factor of $n$ (a notoriously difficult problem) and the property “2|n2|n”, which must be excluded since it occurs for half of integers $n\leqslant x$, but for $\pi(x)-1$ primes $p\leqslant x$ with $n=p+a$ when $a$ is odd, and for only one prime $p$ when $a$ is even.
@@ 1.2. A Kubilius type model for shifted primes
the implied constant in the $\ll$-symbol depending only on $a,A$ and $\alpha$.
@@
The distribution of the large prime factors of shifted primes is not well understood at present. It is known that for infinitely many primes $p$, $P^{+}(p+a)\leqslant p^{0.2844}$ [31] and for infinitely many primes $p$, $P^{+}(p+a)\geqslant p^{0.677}$ [3]. It is conjectured that analogous statements hold with $0.2844$ replaced by any positive number, and with $0.677$ replaced by any number <1<1. We do not address these problems in this paper.
@@ 1.4. An application
for all but $o(x)$ integers up to $x$, as $x\to\infty$. Denote by $q_{1},\ldots,q_{k}$ the primes in $(\xi(x),y]$, let $T_{1}$ be the set of primes in $2,\xi(x)]$, let $T_{j}=\{q_{j-1}\}$ for $2\leqslant j\leqslant k+1$ and set $f_{i}=\omega(n,T_{i})$ for each $i$. The condition (1.8) is then equivalent to $(\omega(n,T_{1}),\ldots,\omega(n,T_{k+1}))\in\mathscr{R}$ for some set $\mathscr{R}\subseteq{\mathbb{N}}_{0}^{k+1}$. Apply Theorem 4 with $\alpha=\frac{1}{2}$ (allowed since Hypothesis $Z(\gamma)$ holds for some $\gamma>\frac{1}{2}$), $A=1$ and $u=\log_{3}x$, and we see that
@@ 3. Sieving by small primes
Then $g_{y}(m)=0$ if $(a,m)>1$ or if $a\equiv m\equiv 1\allowbreak\mkern 7.0mu({\operator@font mod}\,\,2)$, and otherwise equals
@@
Since $\{b\in\mathcal{A}:q|b\}=\pi(x;qm,-a)+O_{a}(1)$ for primes $q$, heuristically we expect that $\Phi_{m}(x,y)$ is approximately $\pi(x)g_{y}(m)$.
@@
If $g_{y}(m)=0$ then either $(a,m)>1$ or $a\equiv m\equiv 1\allowbreak\mkern 7.0mu({\operator@font mod}\,\,2)$, and in this case $\Phi_{m}(x,y)=0$ since $\Phi_{m}(x,y)$ counts only primes $p>|a|+1$. The lemma holds in this case. Also if $y>x^{\theta}/m$ then $u_{m}\leqslant 1$ and, in light of (3.4), the lemma is claiming only that $\Phi_{m}(x,y)\ll\pi(x)/(\phi(m)\log y)$. This bound follows from an upper bound sieve without the need for any prime number estimates.
Now assume $y\leqslant x^{\theta}/m:=D$. When $g_{y}(m)\neq 0$, the lemma follows quickly from the “Fundamental Lemma” of the sieve, e.g., Theorem 2.5’ in [27]. We have
(3.5) $\Phi_{m}(x,y)=\pi(x)g_{y}(m)\Big(1+O_{\varepsilon}(e^{-(1-\varepsilon){u}_{m}^{\phantom{2}}\log{u}_{m}^{\phantom{2}}})\Big)+O_{a}\bigg(D+\sum_{\begin{subarray}{c}d\leqslant D\\ P^{+}(d)\leqslant y\\ (dm,a)=1\end{subarray}}\Big|\pi(x;dm,-a)-\frac{\pi(x)}{\phi(dm)}\Big|\bigg).$
Since $m\not\in\mathcal{E}(x,y;\theta,A+1)$, (3.4) implies that
@@
$\sum_{\begin{subarray}{c}k\geqslant t/d\\ P^{+}(k)\leqslant y\end{subarray}}\frac{1}{k}\leqslant\int_{t/d}^{\infty}\frac{\Psi(w,y)}{w^{2}}\,dw\ll_{\varepsilon}\int_{t/d}^{\infty}w^{-1}e^{-\frac{\log w}{\log y}{\log v}}+\frac{1}{w^{2-\varepsilon}}\,dw\ll(\log y)e^{-v\log v}+(d/t)^{1-\varepsilon}.$
@@ 4. The proof of Theorem 1
$\mathcal{M}_{1}$ $=\{m<x^{\theta},m\not\in\mathcal{E}(x,y;\theta,A+1)\},$
$\mathcal{M}_{2}$ $=\{m<x^{\theta},m\in\mathcal{E}(x,y;\theta,A+1)\},$
$\mathcal{M}_{3}$ $=\{x^{\theta}\leqslant m\leqslant(x+a)/y\},$
$\mathcal{M}_{4}$ $=\{(x+a)/y<m\leqslant x+a\},$
$\mathcal{M}_{5}$ $=\{m>x+a\}.$
Throughout our proof, implied constants may depend on $a$, AA, $\varepsilon$ and $\theta$ only. Now
@@
$\qquad\text{where }B_{\ell}=(\theta u-\ell)\log(\theta u-\ell+1)+\ell\log(\ell+1).$
@@
When $m>(x+a)/y$, $\Phi_{m}(x,y)$ is 1 if mam-a is prime larger than $|a|+1$ and $m-a\leqslant x$, and zero otherwise. Start with the triangle inequality and (3.4), and use the same analysis as in the case $m\in\mathcal{M}_{3}$:
@@
where we applied (3.2) in the last step, assuming that $\delta$ is sufficiently small. By Theorem 1 of [37] [^1] This is stated only for a=1a=-1 but the proof works for any $a$ with no essential changes. and standard bounds for the Dickman function $\rho(u)$ (see, e.g., [40, Ch. III.5]),
$\{|a|+1<p\leqslant x:P^{+}(p+a)\leqslant y\}\ll\pi(x)u\rho(u)\ll\pi(x)e^{-(1-\varepsilon)u\log u}.$